Libby Area Technical Assistance Group, Inc.
PO Box 53 • Libby, MT 59923
November 8th, 2011 • Regular Board Meeting Minutes
(Note: bold items within paragraphs are motions made and voted on as well as action items agreed upon.)
Board Member Attendees:
- Karen Horton
- JoElyn Brus, Treasurer
- Donna Martin, Ad-Hoc
- Tracy McNew, Grant Administrator
- Terry Spear, Technical Advisor via phone
- Carolyn Rutland, MT DEQ
- John Podolinsky, MT DEQ
- David Berry, EPA
- Mel Parker
- Libby Faulk, EPA
- Christina, Progess, EPA
- Rebecca Thomas, EPA
- Liz Fagen, EPA
- Dania Zinner, EPA
- Lorraine Ross, EPA
1. Meeting began at approximately 6:10pm. There was not a quorum.
2. Introductions were made around the room.
3. EPA Report (Rebecca Thomas and Christina Progess)
a. OU1: Excavation of Riverfront Park area and back fill has been completed. Orange fence is to remain in place throughout the winter even when no work is being done. They may do some work at the riverbank in February.
b. OU2: Construction is complete. No post-construction risk assessment yet. A draft of that plan should be available within a month and it will be sent to LATAG.
c. OU3: Surface water sampling took place from April to September but no results are available yet. High levels of asbestos were noted in Rainy Creek so EPA contacted W.R. Grace to discuss instituting a surface water management plan. Keep clean water clean and divert water around known sources of contamination. This plan may also involve Carney and Fleetwood Creeks. Duff has also been collected and shipped to burn site for testing. A site wide plan for sampling bark and duff around NPL boundary will hopefully begin in 2012 and may move on to ABS to determine risk.
d. OU4: Post construction meeting has been done
e. OU5: Waiting for final toxicity value to finalize risk assessment. Will use draft tox values for revising risk assessment but still draft for now.
f. OU6: BNSF submitted complete data to EPA for review. CDM is doing a data gap analysis and once completed EPA will negotiate with them for feasibility study.
4. DEQ Report- John Podolinsky
a. OU7: It was reported that the DEQ office remains open three days per week and that data is being evaluated over the winter.
5. W.R. Grace Report (No Report)
6. Special Presentations
a. Mel Parker requested that EPA review his binder for discussion at a future LATAG meeting.
7. Meeting minutes review and approval (No quorum)
8. Treasurer’s Report (No report)
9. Sub-Committee Reports (No reports)
10. Old Business (None)
11. New Business (None)
12. Open Discussion
a) Mel asked why there was aerial photography being done at the mine site. Christina answered that it can be used to plan surface water management and other remediation. Christina also stated that this photography is being done by Grace and the results are not yet available, but EPA has requested access.
13. The next regular board meeting will be Tuesday December, 6th 2011 at the College.
14. The meeting was ended at approximately 7:00P